# PolyKAN: A Polyhedral Analysis Framework for Provable and Minimal KAN Compression

### v1

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have emerged as a promising alternative to traditional Multi-Layer Perceptrons (MLPs), offering enhanced interpretability and a strong mathematical foundation. However, their parameter efficiency remains a significant challenge for practical deployment. This paper introduces PolyKAN, a novel theoretical framework for KAN compression that provides formal guarantees on both model size reduction and approximation error. By leveraging the inherent piecewise polynomial structure of KANs, we formulate the compression problem as one of optimal polyhedral region merging. We establish a rigorous polyhedral characterization of KANs, develop a complete theory of $\epsilon$-equivalent compression, and design an optimal dynamic programming algorithm that guarantees minimal compression under specified error bounds. Our theoretical analysis demonstrates that PolyKAN achieves provably minimal compression while maintaining strict error control, with polynomial-time complexity in all network parameters. The framework provides the first formal foundation for KAN compression with mathematical guarantees, opening new directions for efficient deployment of interpretable neural architectures.
Keywords: Kolmogorov-Arnold Networks, Model Compression, Polyhedral Analysis, Provable Optimization, Approximation Theory, Computational Complexity
@@ 1 Introduction
The pursuit of interpretable yet powerful neural network architectures has led to renewed interest in mathematically grounded alternatives to traditional black-box models. Kolmogorov-Arnold Networks (KANs) [1], inspired by the celebrated Kolmogorov-Arnold representation theorem [2], have recently emerged as a compelling framework that replaces fixed activation functions with learnable spline transformations. While KANs demonstrate remarkable interpretability and empirical performance, their parameter efficiency presents a substantial obstacle to practical adoption, as each network connection requires an independent spline function with multiple parameters.
Existing approaches to neural network compression—including pruning [3], knowledge distillation [4], and architecture search—largely rely on heuristic strategies without formal guarantees. For KANs specifically, the absence of theoretically grounded compression methods represents a significant gap in the literature. The spline-based foundation of KANs, however, provides a unique opportunity for rigorous analysis through polyhedral theory, as each spline naturally partitions its input domain into piecewise polynomial regions.
In this paper, we present PolyKAN, a comprehensive theoretical framework for KAN compression with provable guarantees. Our work is distinguished by three fundamental contributions:
  1. 1.
    We establish a complete polyhedral characterization of KANs, demonstrating that their input space partitions exhibit a highly regular axis-aligned structure—in stark contrast to the arbitrary polyhedral complexes of ReLU networks. This structural regularity enables rigorous theoretical analysis.
  2. 2.
    We develop a formal theory of $\epsilon$-equivalent compression for KANs, providing necessary and sufficient conditions for region merging while maintaining strict error bounds. Our analysis includes precise error propagation bounds through multiple network layers.
  3. 3.
    We design and analyze an optimal dynamic programming algorithm that guarantees minimal compression under specified error constraints. We prove polynomial-time complexity and establish formal guarantees for both correctness and optimality.
The PolyKAN framework transforms KAN compression from a heuristic process to a mathematically rigorous optimization problem with verifiable guarantees. By bridging spline theory, polyhedral geometry, and algorithm design, we provide a foundation for trustworthy compression of interpretable neural architectures.
@@ 2.1 Kolmogorov-Arnold Networks and Spline Theory
The Kolmogorov-Arnold representation theorem [2] states that any multivariate continuous function can be represented as a finite composition of univariate functions and additions. KANs [1] operationalize this theorem by placing learnable univariate functions (typically B-splines) on network edges while limiting nodes to summation operations. This architectural shift from traditional MLPs provides both mathematical interpretability and empirical advantages in certain function approximation tasks.
@@
where each $s_{j,i}:\mathbb{R}\to\mathbb{R}$ is a spline function parameterized by knot positions and polynomial coefficients.
The approximation properties of splines are well-understood [5], with deep connections to function space theory. Recent work [1] has begun exploring the implications of these properties for neural network design, but a comprehensive theoretical framework for KAN optimization remains underdeveloped.
@@ 2.2 Neural Network Compression
Despite empirical successes, these approaches largely lack formal guarantees. Theoretical work on network compression has typically focused on simplified settings [12] or provided only asymptotic guarantees. The lottery ticket hypothesis [10] offers intriguing insights but falls short of providing constructive compression algorithms with bounded error.
@@ 2.3 Polyhedral Theory in Deep Learning
Polyhedral theory provides powerful tools for analyzing piecewise linear neural networks. A substantial body of work has studied the linear regions of ReLU networks [6, 7, 8], establishing connections between region counts and network expressivity. Mixed-integer programming formulations [9] have been developed for verifying properties of ReLU networks.
However, the polyhedral structure of KANs differs fundamentally from that of ReLU networks. While ReLU networks partition their input space with arbitrarily oriented hyperplanes, KAN partitions are axis-aligned due to their spline-based construction. This structural regularity makes KANs particularly amenable to polyhedral analysis and enables the development of compression algorithms with strong theoretical guarantees.
@@ 3.1 Univariate Spline Partitions
We begin by formalizing the polyhedral structure of individual spline functions, which form the building blocks of KANs.
@@
For a B-spline function $s:[a,b]\to\mathbb{R}$ with knot sequence $t_{0},t_{1},\ldots,t_{k}$ where $a=t_{0}<t_{1}<\cdots<t_{k}=b$, the spline polyhedron is a tuple $\mathcal{P}_{s}=(R,P)$, where:
  • $R=\{R_{1},R_{2},\ldots,R_{k}\}$ is a set of regions with $R_{i}=\{x\in\mathbb{R}:t_{i-1}\leq x\leq t_{i}\}$
  • $P=\{p_{1},p_{2},\ldots,p_{k}\}$ is a set of polynomial functions, with $p_{i}$ defined on $R_{i}$
@@
3.2 Multi-Layer KAN Partitions
For multi-layer KANs, the polyhedral structure emerges from the composition of spline functions across layers. Consider an $L$-layer KAN with architecture $n_{0},n_{1},\ldots,n_{L}]$, where $n_{l}$ denotes the number of nodes in layer $l$.
@@ 3.2 Multi-Layer KAN Partitions
The input space partition of an $L$-layer KAN is the refinement of partitions induced by spline functions across all layers. The total number of linear regions satisfies:
@@
where $k^{(l)}_{j,i}$ is the number of knots in spline function $s^{(l)}_{j,i}$ connecting node $i$ in layer l1l-1 to node $j$ in layer $l$.
@@
Each spline function $s^{(l)}_{j,i}$ partitions its one-dimensional input space into at most $(k^{(l)}_{j,i}-1)$ intervals. Since these partitions act independently along different dimensions at each layer, the Cartesian product of these partitions yields the region structure for that layer’s output space. The composition across layers refines these partitions, leading to the product bound. ∎
@@ 3.3 Structural Properties of KAN Polyhedra
Lemma 1 (Axis-Alignment).
@@
The knot points of each spline function $s^{(l)}_{j,i}(x_{i})$ define partitions of the form $x_{i}=t_{m}$, which are hyperplanes orthogonal to the $i$-th coordinate axis. The composition of such axis-aligned partitions preserves axis-alignment. ∎
@@
Every KAN polyhedral region is a axis-aligned rectangle (the Cartesian product of intervals).
@@
Following from the axis-alignment property, the intersection of half-spaces defined by axis-aligned hyperplanes necessarily produces rectangular regions. ∎
@@
Within each polyhedral region, a KAN is a smooth multivariate polynomial function. Across region boundaries, the KAN maintains continuity (for linear B-splines) or higher-order smoothness (for higher-degree B-splines).
These structural properties significantly simplify polyhedral analysis compared to ReLU networks, where regions can be arbitrary convex polyhedra and boundaries have arbitrary orientations.
4 Provable KAN Compression Theory
@@ 4.1 Formal Problem Statement
We now formalize the KAN compression problem with provable guarantees.
@@
Given a KAN network $\mathcal{N}:\mathcal{X}\to\mathbb{R}$ and error tolerance $\epsilon>0$, a compressed network $\mathcal{N}^{\prime}$ is an $\epsilon$-equivalent compression of $\mathcal{N}$ if:
@@
Our objective is to find an $\epsilon$-equivalent compression $\mathcal{N}^{\prime}$ with minimal number of parameters (or equivalently, minimal total knots across all spline functions).
4.2 Region Mergability Theory
The foundation of our compression approach is the theory of merging adjacent polyhedral regions while controlling approximation error.
@@ 4.2 Region Mergability Theory
Let $R_{i}$ and $R_{j}$ be two adjacent polyhedral regions in a KAN, with corresponding polynomial functions $p_{i}$ and $p_{j}$. The regions are $\epsilon$-mergable if there exists a single polynomial $p_{ij}$ such that:
@@
This definition captures the intuition that we can replace two different polynomials defined on adjacent regions with a single polynomial that approximates both within tolerance $\epsilon$.
@@
Let $t_{m}$ be an interior knot of spline function $s(x)$, with adjacent regions $R_{m-1}=[t_{m-1},t_{m}]$ and $R_{m}=[t_{m},t_{m+1}]$. If $R_{m-1}$ and $R_{m}$ are $\epsilon$-mergable, then knot $t_{m}$ can be eliminated while maintaining $\epsilon$-equivalence.
@@
If $R_{m-1}$ and $R_{m}$ are $\epsilon$-mergable, there exists a polynomial $p$ such that $|s(x)-p(x)|\leq\epsilon$ for all $x\in R_{m-1}\cup R_{m}$. By replacing the original piecewise representation with $p$ on the merged region $t_{m-1},t_{m+1}]$, we obtain a new spline function $\tilde{s}$ that satisfies $\|s-\tilde{s}\|_{\infty}\leq\epsilon$ while eliminating one knot. ∎
@@ 4.3 Error Propagation Analysis
For multi-layer KAN compression, we must understand how compression errors propagate through the network.
@@
Consider a KAN layer $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ with:
@@
Taking the supremum over all $\mathbf{x}$ and maximum over $j$ completes the proof. ∎
This lemma enables hierarchical error budgeting across network layers. Given a global error tolerance $\epsilon$, we can allocate error budgets $\delta_{l}$ to each layer $l$ such that the cumulative effect satisfies the global bound.
4.4 Optimal Compression Formulation
We now formulate KAN compression as a combinatorial optimization problem. Let $z_{l,i,j,m}\in\{0,1\}$ indicate whether the $m$-th knot of spline $s^{(l)}_{j,i}$ is retained, and let $\delta_{l,i,j}$ be the error budget allocated to this spline.
The optimal compression problem becomes:
$\min$ $\sum_{l=1}^{L}\sum_{i=1}^{n_{l-1}}\sum_{j=1}^{n_{l}}\sum_{m=1}^{k_{l,i,j}}z_{l,i,j,m}$ (5)
s.t. $\epsilon$-equivalence constraints
Knot sequence feasibility
$z_{l,i,j,m}\in\{0,1\}$
While this integer programming formulation captures the optimality structure, its direct solution is computationally challenging. In the next section, we develop an efficient dynamic programming algorithm that exploits the problem structure to achieve optimality with polynomial complexity.
5 The PolyKAN Compression Algorithm
5.1 Algorithm Overview
We present the PolyKAN compression algorithm, which systematically reduces KAN complexity while guaranteeing $\epsilon$-equivalence. The algorithm operates layer by layer, compressing each spline function optimally using dynamic programming.
Algorithm 1 PolyKAN Provable Compression Algorithm
1:Original KAN network $\mathcal{N}$, global error bound $\epsilon$
2:Compressed KAN network $\mathcal{N}^{\prime}$ with $\|\mathcal{N}-\mathcal{N}^{\prime}\|_{\infty}\leq\epsilon$
[…]
4:for $l\leftarrow 1$ to $L$ do
5:  for each spline function $s$ in layer $l$ do
6:   $\tilde{s}\leftarrow\text{OptimalSplineCompression}(s,\delta_{l})$ $\triangleright$ Algorithm 2

### v2

# PolyKAN: A Polyhedral Analysis Framework for Provable and Approximately Optimal KAN Compression

@@ Abstract
Kolmogorov-Arnold Networks (KANs) have emerged as a promising alternative to traditional Multi-Layer Perceptrons (MLPs), offering enhanced interpretability and a solid mathematical foundation. However, their parameter efficiency remains a significant challenge for practical deployment. This paper introduces PolyKAN, a novel theoretical framework for KAN compression that provides formal guarantees on both model size reduction and approximation error. By leveraging the inherent piecewise polynomial structure of KANs, we formulate the compression problem as a polyhedral region merging task. We establish a rigorous polyhedral characterization of KANs, develop a complete theory of $\epsilon$-equivalent compression, and design a dynamic programming algorithm that achieves approximately optimal compression under specified error bounds. Our theoretical analysis demonstrates that PolyKAN achieves provably near-optimal compression while maintaining strict error control, with guaranteed global optimality for univariate spline functions. This framework provides the first formal foundation for KAN compression with mathematical guarantees, opening new directions for the efficient deployment of interpretable neural architectures.
Keywords: Kolmogorov-Arnold Networks, Model Compression, Polyhedral Analysis, Provable Optimization, Approximation Theory, Dynamic Programming
@@ 1 Introduction
The pursuit of neural network architectures that combine interpretability with strong performance has renewed interest in mathematically grounded alternatives to traditional black-box models. Kolmogorov-Arnold Networks (KANs) [1], inspired by the celebrated Kolmogorov-Arnold Representation Theorem [2], have recently emerged as a compelling framework that replaces fixed activation functions with learnable spline transformations. Although KANs demonstrate remarkable interpretability and empirical performance in certain function approximation tasks, their parameter efficiency presents a major obstacle to practical application, as each network connection requires an independent spline function with multiple parameters.
Existing neural network compression methods—including pruning [3], knowledge distillation [4], and architecture search—largely rely on heuristic strategies and lack formal guarantees. For KANs in particular, the absence of theoretically grounded compression methods represents a significant gap in the literature. However, the spline-based roots of KANs offer unique opportunities for rigorous analysis through polyhedral theory, as each spline naturally partitions its input domain into piecewise polynomial regions.
This paper presents PolyKAN, a comprehensive theoretical framework for KAN compression with provable guarantees. Our work makes three fundamental contributions: First, we establish a complete polyhedral characterization of KANs, demonstrating that their input space partitions exhibit highly regular axis-aligned structures—in stark contrast to the arbitrary polyhedral complexes of ReLU networks. This structural regularity enables rigorous theoretical analysis. Second, we develop a formal theory of $\epsilon$-equivalent compression for KANs, providing necessary and sufficient conditions for region merging while preserving strict error bounds. Our analysis includes precise bounds on error propagation through multiple network layers. Third, we design and analyze a dynamic programming algorithm that guarantees optimal compression for univariate spline functions and delivers approximately optimal guarantees for the entire network, with provable polynomial time complexity.
The PolyKAN framework transforms KAN compression from a heuristic process into a mathematically rigorous optimization problem with verifiable guarantees. By connecting spline theory, polyhedral geometry, and algorithm design, we lay the foundation for trustworthy compression of interpretable neural architectures.
@@ 2.1 Kolmogorov-Arnold Networks and Spline Theory
The Kolmogorov-Arnold Representation Theorem [2] states that any multivariate continuous function can be represented as a composition of a finite number of univariate functions and additions. KANs [1] implement this theorem by placing learnable univariate functions (typically B-splines) on the network edges while restricting nodes to summation operations. This architectural shift from traditional MLPs provides advantages in both mathematical interpretability and empirical performance in certain function approximation tasks.
@@
where each $s_{j,i}:\mathbb{R}\to\mathbb{R}$ is a spline function parameterized by knot locations and polynomial coefficients.
The approximation properties of splines are well-understood [5] and have deep connections to function space theory. Recent work [1] has begun to explore the implications of these properties for neural network design, but a comprehensive theoretical framework for KAN optimization remains underdeveloped.
@@ 2.2 Neural Network Compression
Despite empirical success, most of these methods lack formal guarantees. Theoretical work on network compression typically focuses on simplified settings [12] or provides only asymptotic guarantees. The lottery ticket hypothesis [10] offers intriguing insights but fails to provide constructive compression algorithms with bounded error.
@@ 2.3 Polyhedral Theory in Deep Learning
Polyhedral theory provides powerful tools for analyzing piecewise linear neural networks. Substantial work has studied the linear regions of ReLU networks [6, 7, 8], establishing connections between region counts and network expressive power. Mixed-integer programming formulations [9] have been developed for verifying properties of ReLU networks.
Recent work has begun to explore connections between KANs and ReLU networks. [13] established formal relationships between piecewise linear KANs and ReLU networks, proving that under certain conditions, both architectures exhibit similar polyhedral complexity. Meanwhile, [14] introduced ReLU-KANs, a variant that replaces spline functions with ReLU activations while maintaining the Kolmogorov-Arnold structure, further bridging the gap between these architectures.
However, the polyhedral structure of general KANs differs fundamentally from that of ReLU networks. While ReLU networks partition their input space with hyperplanes of arbitrary orientation, KAN partitions are axis-aligned due to their spline-based construction. This structural regularity makes KANs particularly amenable to polyhedral analysis and enables the development of compression algorithms with strong theoretical guarantees.
@@ 3.1 Univariate Spline Partitions
We begin by formalizing the polyhedral structure of individual spline functions, which constitute the building blocks of KANs.
@@
For a B-spline function $s:[a,b]\to\mathbb{R}$ defined on interval $[a,b]$ with knot sequence $t_{0},t_{1},\ldots,t_{k}$ where $a=t_{0}<t_{1}<\cdots<t_{k}=b$, its spline polyhedron is a tuple $\mathcal{P}_{s}=(R,P)$ where:
  • $R=\{R_{1},R_{2},\ldots,R_{k}\}$ is a collection of regions with $R_{i}=\{x\in\mathbb{R}:t_{i-1}\leq x\leq t_{i}\}$
  • $P=\{p_{1},p_{2},\ldots,p_{k}\}$ is a collection of polynomial functions with $p_{i}$ defined on $R_{i}$
@@
3.2 Multilayer KAN Partitions
For multilayer KANs, the polyhedral structure arises from the composition of spline functions across layers. Consider an $L$-layer KAN with architecture $n_{0},n_{1},\ldots,n_{L}]$, where $n_{l}$ denotes the number of nodes in layer $l$.
@@ 3.2 Multilayer KAN Partitions
The input space partition of an $L$-layer KAN is a refinement of the partitions induced by spline functions across all layers. The total number of linear regions satisfies:
@@
where $k^{(l)}_{j,i}$ is the number of knots in the spline function $s^{(l)}_{j,i}$ connecting node $i$ in layer l1l-1 to node $j$ in layer $l$.
@@
Each spline function $s^{(l)}_{j,i}$ partitions its one-dimensional input space into at most $(k^{(l)}_{j,i}-1)$ intervals. Since these partitions act independently along different dimensions at each layer, the Cartesian product of these partitions produces the region structure of that layer’s output space. Composition across layers refines these partitions, leading to the product bound. ∎
@@ 3.3 Structural Properties of KAN Polyhedra
Lemma 1 (Axis-Alignedness).
@@
The knots of each spline function $s^{(l)}_{j,i}(x_{i})$ define partitions of the form $x_{i}=t_{m}$, which are hyperplanes perpendicular to the $i$-th coordinate axis. The composition of such axis-aligned partitions preserves axis-alignedness. ∎
@@
Each KAN polyhedral region is an axis-aligned rectangle (Cartesian product of intervals).
@@
By the axis-aligned property, intersections of half-spaces defined by axis-aligned hyperplanes necessarily produce rectangular regions. ∎
@@
Within each polyhedral region, a KAN is a smooth multivariate polynomial function. Across region boundaries, KANs maintain continuity (for linear B-splines) or higher-order smoothness (for higher-degree B-splines).
These structural properties significantly simplify polyhedral analysis compared to ReLU networks, whose regions can be arbitrary convex polyhedra with boundaries of arbitrary orientation.
4 Theory of Provable KAN Compression
@@ 4.1 Formal Problem Statement
We now formalize the problem of KAN compression with provable guarantees.
@@
Given a KAN network $\mathcal{N}:\mathcal{X}\to\mathbb{R}$ and an error tolerance $\epsilon>0$, a compressed network $\mathcal{N}^{\prime}$ is an $\epsilon$-equivalent compression of $\mathcal{N}$ if:
@@
Problem 1 (Optimal KAN Compression).
Given a KAN network $\mathcal{N}$ and an error tolerance $\epsilon>0$, find an $\epsilon$-equivalent compression $\mathcal{N}^{\prime}$ that minimizes the total number of knots across all spline functions.
4.2 Computational Complexity Analysis
We first analyze the computational complexity of the optimal KAN compression problem.
Theorem 2 (NP-Hardness).
The optimal KAN compression problem is NP-hard.
Proof.
Consider a simplified version: given a set of one-dimensional intervals $I_{1},I_{2},\ldots,I_{n}$ with corresponding polynomials $p_{1},p_{2},\ldots,p_{n}$, and an error tolerance $\epsilon$, find the smallest set of knots such that each merged interval can be approximated by a single polynomial with error $\leq\epsilon$. This can be reduced to the set cover problem, which is classically NP-complete.
More specifically, given a set cover instance $(U,\mathcal{S})$ where $U$ is the universe and $\mathcal{S}$ is a family of subsets, we can construct a KAN compression instance by mapping each element in $U$ to an interval and each subset in $\mathcal{S}$ to a feasible interval merging. The optimal compression corresponds to the minimum set cover. ∎
Although the overall problem is NP-hard, the special structure of KANs allows us to design efficient approximation algorithms.
4.3 Theory of Region Mergability
The foundation of our compression approach is a theory of merging adjacent polyhedral regions while controlling approximation error.
@@ 4.3 Theory of Region Mergability
Let $R_{i}$ and $R_{j}$ be two adjacent polyhedral regions in a KAN with corresponding polynomial functions $p_{i}$ and $p_{j}$. These regions are $\epsilon$-mergable if there exists a single polynomial $p_{ij}$ such that:
@@
This definition captures the intuition that we can replace two different polynomials defined on adjacent regions with a single polynomial that approximates both original polynomials within tolerance $\epsilon$.
@@
Let $t_{m}$ be an interior knot of a spline function $s(x)$ with adjacent regions $R_{m-1}=[t_{m-1},t_{m}]$ and $R_{m}=[t_{m},t_{m+1}]$. If $R_{m-1}$ and $R_{m}$ are $\epsilon$-mergable, then knot $t_{m}$ can be eliminated while preserving $\epsilon$-equivalence.
@@
If $R_{m-1}$ and $R_{m}$ are $\epsilon$-mergable, there exists a polynomial $p$ such that $|s(x)-p(x)|\leq\epsilon$ for all $x\in R_{m-1}\cup R_{m}$. By replacing the original piecewise representation on the merged region $t_{m-1},t_{m+1}]$ with $p$, we obtain a new spline function $\tilde{s}$ satisfying $\|s-\tilde{s}\|_{\infty}\leq\epsilon$ while eliminating one knot. ∎
@@ 4.4 Error Propagation Analysis
For multilayer KAN compression, we must understand how compression errors propagate through the network.
@@
Consider a KAN layer $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{m}$ where:
@@
Taking the supremum over all $\mathbf{x}$ and the maximum over $j$ completes the proof. ∎
This lemma enables layered error budget allocation across network layers. Given a global error tolerance $\epsilon$, we can allocate error budgets $\delta_{l}$ to each layer $l$ such that the cumulative effect satisfies the global bound.
5 Approximately Optimal Compression Algorithms
5.1 Optimal Compression for Univariate Splines
Although overall KAN compression is NP-hard, for univariate spline functions we can design efficient algorithms with optimality guarantees.
Algorithm 1 Optimal Compression for Single Spline
1:Spline function $s$, knot sequence $t_{0},t_{1},\ldots,t_{k}$, error tolerance $\epsilon$
2:Compressed knot sequence
3:Initialize DP table: $dp[i]\leftarrow i+1$ for $i=0,1,\ldots,k$ $\triangleright$ $dp[i]$: min knots to reach $t_{i}$
4:Initialize backtrack pointers: $prev[i]\leftarrow-1$ for $i=0,1,\ldots,k$
5:for $i=1$ to $k$ do
6:  for $j=0$ to i1i-1 do
7:   Check if interval $t_{j},t_{i}]$ can be approximated by a single polynomial $p$ with error $\leq\epsilon$
8:   if mergable and $dp[j]+1<dp[i]$ then
9:     $dp[i]\leftarrow dp[j]+1$
10:     $prev[i]\leftarrow j$
11:   end if
12:  end for
13:end for
14:return Optimal knot sequence constructed by backtracking prevprev array
Theorem 4 (Univariate Optimality).
Algorithm 1 guarantees the global optimal solution for the single spline optimal compression problem.
Proof.
The algorithm is a classic interval partitioning dynamic programming. The optimal substructure property holds: the optimal solution from $t_{0}$ to $t_{i}$ must consist of the optimal solution from $t_{0}$ to some $t_{j}$ plus a single polynomial approximation for interval $t_{j},t_{i}]$. The dynamic programming correctly explores all possible partition points. ∎
Theorem 5 (Time Complexity).
Algorithm 1 has time complexity $O(k^{3}\cdot T_{\text{fit}})$, where $k$ is the original number of knots and $T_{\text{fit}}$ is the time for polynomial fitting.
Proof.
The algorithm has $O(k^{2})$ state transitions, each requiring $O(k)$ time to check the feasibility of interval merging (via polynomial fitting and error computation), yielding total complexity $O(k^{3}\cdot T_{\text{fit}})$. ∎
5.2 Approximate Compression for Multilayer KANs
For entire KAN networks, we employ a layered compression strategy.
Algorithm 2 Approximate Compression for Multilayer KAN
1:KAN network $\mathcal{N}$, global error tolerance $\epsilon$
2:Compressed KAN network $\mathcal{N}^{\prime}$
3:Allocate global error budget $\epsilon$ proportionally across layers: $\epsilon=\sum_{l=1}^{L}\epsilon_{l}$
4:for each layer $l=1$ to $L$ do
5:  for each spline function $s$ in this layer do
6:   Compress spline $s$ using Algorithm 1 with error budget $\epsilon_{l}/n_{l}$